# and Bethe Center for Theoretical Physics, Universität Bonn, 53115 Bonn, Germany

### v2

@@ 1 Introduction
Obviously, the condition $R/L\ll 1$ is violated, when the scattering in the presence of the electromagnetic interactions is considered (QCD+photons on the lattice). Moreover, at physical quark masses the pions are rather light, which leads to problems in the study of nucleon-nucleon interactions on the lattice. Namely, as explicitly demonstrated in the recent paper Meng:2021uhz , the partial-wave mixing at the physical point in the Lüscher equation for the $NN$ scattering is indeed substantial. A closely related problem is the appearance of the so-called $t$-channel (left-hand) cut in the $NN$ scattering amplitude running from negative infinity till $s=(2m_{N})^{2}-M_{\pi}^{2}$ along the real axis in the complex $s$-plane, see, e.g., Raposo:2023nex ; Raposo:2023oru (Here, $m_{N}$ and $M_{\pi}$ denote the nucleon and the pion masses, respectively.). Since $M_{\pi}\ll m_{N}$ in nature, the gap between the $t$-channel cut and the right-hand unitarity cut is very small. On the other hand, as seen in the latest studies of the $NN$ scattering on the lattice, many energy levels are located on the $t$-channel cut (see, e.g., Fig. 3 in Ref. Green:2021qol ). The standard Lüscher approach is obviously not applicable in this case. Note also that $NN$ scattering is not only the physically interesting process where these problems emerge. For example, in the description of the $T_{cc}(3875)^{+}$ state the same problem shows up in full glory. Moreover, it can be seen that the structure of the singularities is completely different in the two cases $M_{\pi}<M_{D^{*}}-M_{D}$ and $M_{\pi}>M_{D^{*}}-M_{D}$, and hence reveals a critical dependence on the values of quark masses used in the lattice calculations Du:2023hlu .
@@
The aim of the present paper is to address the problem of the long-range force in the Lüscher equation in a general fashion and to derive a modified Lüscher equation, which has a much larger domain of applicability than the original one. To simplify life as much as possible, in this paper we do not consider the theories with massless particles – the inclusion of QED is relegated to future publications. Furthermore, we ignore purely technical issues like the inclusion of spin, relativistic kinematics or moving frames. The key observation that allows one to achieve the stated objective is that the long-range part of the potential, which gives rise to all above problems, is usually well known and can be expressed in terms of few parameters that can be accurately measured on the lattice (like the axial-vector coupling constant $g_{A}$ or the pion decay constant $F_{\pi}$ in case of the one-pion exchange). The short-range part of potential is unknown and should be fitted to the lattice data on the two-body energy levels by using the modified Lüscher equation.
@@ 2.1 Modified effective range expansion
Finally, in the splitting $V(r)=V_{L}(r)+V_{S}(r)$, one could modify both $V_{L}(r)$ and $V_{S}(r)$, adding and subtracting the same string of the short-range Yukawa terms. This will render $V_{L}(r)$ superregular and will not change the interpretation of $V_{S}(r)$ as a short-range potential.
@@ 2.2 NREFT framework
The first line gives the Lippmann-Schwinger for $T_{L}$, while the second line gives the expression for $G_{L}$ in terms of $G_{0}$ and $T_{L}$. Both quantities are needed for the NREFT formulation of the modified effective range expansion. The explicit expressions read
@@
Now, let us consider Born series for the quantity
@@ 2.3 Non-derivative interactions
where $M_{\ell}(q_{0})$ is given by Eq. (4). The real polynomial can be safely dropped as it corresponds to the choice of the renormalization prescription. Next, we require the scattering wave function $\phi_{0}(k,r)\doteq\phi_{0}(k,k,r)$ defined in Eq. (86). Using
@@ 2.4 Derivative interactions
Consider now the situation when the matrix element of the potential $V_{S}$ is a generic low-energy polynomial defined in Eq. (10). This is no more true for the potential $\tilde{V}_{S}({\bf p},{\bf q})$, defined in Eq. (27). Here, we wish to address the structure of the latter in more detail. The partial-wave expansion of the $V_{S}$ has the form:
Convoluting Eq. (8) with the wave functions, integrals of the following type emerge
@@
Finally, using Eqs. (28) and (35), one arrives at the modified effective-range expansion as given in Eq. (3), with $K_{\ell}^{M}(q_{0}^{2})$ being by a low-energy polynomial.
@@ 3.1 Derivation of the quantization condition
Furthermore, the finite-volume spectrum of the system is determined by the pole positions of the full $T$ matrix that can be written down in a form of a finite-volume analog of Eq. (17). Using the fact that the poles of $T_{L}$ will eventually cancel Doring:2009bi (see also Appendix B), it is straightforward to conclude that the spectrum will be determined by the poles of $T_{S}$. Moreover, it is easily seen that no spurious poles emerge, since the poles that emerge in $G_{L}$ are shifted by the short-range interaction. Using now the basis of eigenfunctions of the Hamiltonian $H_{L}$ and defining the quantity
@@
The next steps in the derivation repeat those in the infinite volume. We use following ansatz for the matrix $B$
@@ 3.2.1 Negative energies, $q_{0}^{2}<0$
$H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})=(H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})-H^{\infty}_{\ell m,\ell^{\prime}m^{\prime}}(q_{0}))+4\pi\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}M_{\ell}(q_{0})\,.$ (66)
@@
The first term here is the standard Lüscher zeta-function. In order to calculate the remaining two terms, let us consider the Lippmann-Schwinger equation for $T_{L}(q_{0}^{2})$ (a finite-volume counterpart of Eq. (2.2)). Carrying out the partial-wave expansion

### v3

@@ 1 Introduction
Obviously, the condition $R/L\ll 1$ is violated, when the scattering in the presence of the electromagnetic interactions is considered (QCD+photons on the lattice). Moreover, at physical quark masses the pions are rather light, which leads to problems in the study of nucleon-nucleon interactions on the lattice. Namely, as explicitly demonstrated in the recent paper Meng:2021uhz , the partial-wave mixing at the physical point in the Lüscher equation for the $NN$ scattering is indeed substantial. A closely related problem is the appearance of the so-called $t$-channel (left-hand) cut in the $NN$ partial-wave scattering amplitudes, running from negative infinity till $s=(2m_{N})^{2}-M_{\pi}^{2}$ along the real axis in the complex $s$-plane, see, e.g., Raposo:2023nex ; Raposo:2023oru (Here, $m_{N}$ and $M_{\pi}$ denote the nucleon and the pion masses, respectively.). Since $M_{\pi}\ll m_{N}$ in nature, the gap between the $t$-channel cut and the right-hand unitarity cut is very small. On the other hand, as seen in the latest studies of the $NN$ scattering on the lattice, many energy levels are located on the $t$-channel cut (see, e.g., Fig. 3 in Ref. Green:2021qol ). The standard Lüscher approach is obviously not applicable in this case. Note also that $NN$ scattering is not only the physically interesting process where these problems emerge. For example, in the description of the $T_{cc}(3875)^{+}$ state the same problem shows up in full glory. Moreover, it can be seen that the structure of the singularities is completely different in the two cases $M_{\pi}<M_{D^{*}}-M_{D}$ and $M_{\pi}>M_{D^{*}}-M_{D}$, and hence reveals a critical dependence on the values of quark masses used in the lattice calculations Du:2023hlu .
@@
The aim of the present paper is to address the problem of the long-range force in the Lüscher equation in a general fashion and to derive a modified Lüscher equation, which has a much larger domain of applicability than the original one. To simplify life as much as possible, in this paper we do not consider the theories with massless particles – the inclusion of QED is relegated to future publications. Furthermore, we ignore purely technical issues like the inclusion of spin, relativistic kinematics or moving frames. The key observation that allows one to achieve the stated objective is that the long-range part of the potential, which gives rise to all above problems, is usually well known and can be expressed in terms of few parameters that can be accurately measured on the lattice. The short-range part of potential is unknown and should be fitted to the lattice data on the two-body energy levels by using the modified Lüscher equation.
@@ 2.1 Modified effective range expansion
Finally, in the splitting $V(r)=V_{L}(r)+V_{S}(r)$, one could modify both $V_{L}(r)$ and $V_{S}(r)$, adding and subtracting the same string of the short-range Yukawa terms. This will render $V_{L}(r)$ superregular and will not change the interpretation of $V_{S}(r)$ as a short-range potential. Hence, the above regularization serves solely the purpose of recasting the potential into the form that obeys the requirements of Ref. vanHaeringen:1981pb .
@@ 2.2 NREFT framework
The first line gives the LS equation for $T_{L}$, while the second line gives the expression for $G_{L}$ in terms of $G_{0}$ and $T_{L}$. Both quantities are needed for the NREFT formulation of the modified effective range expansion. The explicit expressions read
@@
Now, let us consider the Born series for the quantity
@@ 2.3 Non-derivative interactions
where $M_{\ell}(q_{0})$ is given by Eq. (4). The real polynomial can be safely dropped as it corresponds to the choice of the renormalization prescription.
Next, we express the wave function at the origin, which appears in Eq. (31), through the scattering wave function $\phi_{0}(k,r)\doteq\phi_{0}(k,k,r)$ defined in Eq. (87). Using
@@ 2.4 Derivative interactions
Consider now the situation when the matrix element of the potential $V_{S}$ is a generic low-energy polynomial defined in Eq. (10). This is no more true for the potential $\tilde{V}_{S}({\bf p},{\bf q})$, defined in Eq. (27). Here, we wish to address the structure of the latter in more detail. The partial-wave expansion of the $V_{S}$ is given in Eq. (8). Convoluting this equation with the wave functions, integrals of the following type emerge
@@
Finally, using Eqs. (28) and (35), one arrives at the modified effective-range expansion as given in Eq. (3), with $K_{\ell}^{M}(q_{0}^{2})$ being a low-energy polynomial.
@@ 3.1 Derivation of the quantization condition
Furthermore, the finite-volume spectrum of the system is determined by the pole positions of the full $T$-matrix that can be written down in a form of a finite-volume analog of Eq. (17). Using the fact that the poles of $T_{L}$ will eventually cancel Doring:2009bi (see also Appendix B), it is straightforward to conclude that the spectrum will be determined by the poles of $T_{S}$. Moreover, it is easily seen that no spurious poles emerge, since the poles that emerge in $G_{L}$ are shifted by the short-range interaction. Using now the basis of eigenfunctions of the Hamiltonian $H_{L}$ and defining the quantity
@@
The next steps in the derivation repeat those in the infinite volume. We use the following ansatz for the matrix $B$
@@ 3.2.1 Negative energies, $q_{0}^{2}<0$
$H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})=(H_{\ell m,\ell^{\prime}m^{\prime}}(q_{0})-H^{\infty}_{\ell m,\ell^{\prime}m^{\prime}}(q_{0}))+\frac{1}{4\pi}\,\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}M_{\ell}(q_{0})\,.$ (66)
@@
The first term here is the standard Lüscher zeta-function. In order to calculate the remaining two terms, let us consider the LS equation for $T_{L}(q_{0}^{2})$ (a finite-volume counterpart of Eq. (2.2)). Carrying out the partial-wave expansion
@@ 3.4 Exponentially suppressed effects
3.5 The case with the short-range potential only
The limit $V_{L}(r)\to 0$ is trivial. In this limit, the Jost function $f_{L}(q)=1$, $\sigma_{\ell}(q)=0$, and Eq. (3) reduces to the familiar expression
$K_{\ell}^{M}(q^{2})=q^{2\ell+1}\cot\delta_{\ell}(q)\,.$ (84)
Furthermore, the modified Lüscher zeta-function $H_{\ell m,\ell^{\prime}m^{\prime}}$ reduces to the conventional one, $Z_{\ell m,\ell^{\prime}m^{\prime}}$, and Eq. (65) turns into the original Lüscher equation.